# Multimodal Dataset Distillation Made Simple by Prototype-guided Data Synthesis

### v1

@@ 1 Introduction
A variety of approaches have been proposed to reduce dataset size, including coreset selection (Farahani and Hekmatfar, 2009; Welling, 2009), dataset filtering (Gadre et al., 2023), and pruning methods (Yang et al., 2023; Mahmoud et al., 2024), which have shown competitive performance in the context of CLIP. However, since they rely on representative subsets of the original dataset, they are effective only with sufficiently large reduced datasets and fail once the subset is too small to preserve semantic diversity. In contrast, dataset distillation (Wang et al., 2018) synthesizes new samples by interpolating the semantics encoded in the original dataset, producing compact yet highly informative datasets. This makes it particularly effective on extremely reduced datasets, where it has outperformed selection-based strategies in image classification (Zhao and Bilen, 2021; Cazenavette et al., 2022), even with only a few distilled samples per class.
@@ 3.1 Problem formulation
where $\mathcal{L}$ denotes a contrastive loss such as InfoNCE. In practice, since text is inherently discrete, dataset distillation is performed in the text embedding space, while images are parameterized in the pixel space. Existing methods typically rely on optimization-based approaches, which require repeated training on the full dataset and storage of intermediate parameters, resulting in significant computational and storage overhead. Moreover, their performance is architecture-dependent, often degrading when used to train models with different backbones.
@@ 3.2 Our PDS framework
Cluster matching for prototype construction. While we obtain semantically coherent clusters within each modality, semantically related clusters are not aligned across modalities. To establish cross-modal correspondences, we measure the association between $C_{i}^{\text{img}}$ and $C_{j}^{\text{txt}}$ by the number of shared image-text pairs whose embeddings belong to the respective clusters. Clusters with more shared pairs are regarded as more semantically related, and we therefore aim to find a one-to-one cluster matching that maximizes the total number of shared pairs.
@@ 4.1 Experimental Settings
Implementation details. Unless otherwise specified, all baselines use ViT-L/14 CLIP model (Radford et al., 2021). For PDS, images are generated with the unCLIP decoder using classifier-free guidance (Ho and Salimans, 2022) with a guidance scale of 5.0 and 100 sampling steps, and resized to $224\times 224$ for evaluation. To assess cross-architecture generalization, we follow Xu et al. (2024) and evaluate whether the distilled datasets remain effective when transferred to alternative vision backbones such as ResNet-50 (He et al., 2016) and ViT-Ti/16 (Dosovitskiy et al., 2021). For this evaluation, we use each alternative backbone as the image encoder and attach a learnable linear projection layer after the text encoder. We then fine-tune the CLIP-style model on the distilled dataset while keeping the text encoder frozen. For all baselines, we adopt the hyperparameters reported in the original papers, and conduct all experiments on a single RTX 3090 GPU.
@@ 4.2 Performance Comparison
We evaluate PDS against three categories of baselines: multimodal dataset distillation methods, dataset subset selection methods, and learning-free distillation methods for image classification. These comparisons are designed to highlight three aspects of our contribution: (i) mitigating the cross-generalization limitations of existing multimodal dataset distillation methods, (ii) demonstrating the advantage of dataset distillation over subset selection when the reduced dataset is extremely small, and (iii) overcoming the limitations of existing learning-free distillation approaches when extending from image classification to multimodal datasets. Detailed descriptions of each baseline are provided in Appendix A.
@@
Dataset subset selection. To obtain a highly reduced yet informative dataset, representative images must preserve broad semantic coverage. In this case, dataset distillation offers a more effective solution than subset selection, as it synthesizes data by interpolating diverse semantic features. To validate this, we evaluate PDS against classical coreset-selection methods, K-center (Farahani and Hekmatfar, 2009) and Herding (Welling, 2009), as well as widely used filtering schemes such as CLIP score filtering, LAION filtering, and image-based filtering (Gadre et al., 2023).
As shown in Table 2, PDS consistently achieves the highest performance, outperforming the strongest baseline (Herding) by 17.2 and 10.8 $\mathrm{pp}$ in IR@10 and TR@10, respectively. Although dataset filtering methods are widely used in CLIP training due to their effectiveness with moderately reduced datasets, they fail on extremely reduced datasets. These results demonstrate that dataset distillation is particularly well-suited for such scenarios, as it preserves broader semantic coverage by interpolating diverse semantic features and achieves superior performance.
@@ 4.3 Ablation studies
Advantage of using a generative model. We analyze the benefits of using a generative model by comparing PDS with optimization-based multimodal dataset distillation methods. Such optimization-based approaches often synthesize images nearly identical to the initialization sample (Figure 3, left), essentially adding architecture-dependent adversarial perturbations rather than generating genuinely new data. In contrast, PDS employs a generative model to synthesize images from prototypes constructed by interpolating diverse semantic features.
@@
Impact of image prototype on image generation. To assess the impact of incorporating an image prototype into image generation, we compare PDS with three baselines that obtain images through different mechanisms. First, we consider two baselines that do not use image prototypes. The first is text-prototype-based retrieval, which retrieves the caption in the training set whose embedding has the highest cosine similarity to the given text prototype and then selects its paired real image. The second is an unCLIP-based text-to-image generation method (Ramesh et al., 2022), which applies the full unCLIP pipeline (prior and decoder) to generate an image from the retrieved caption. In addition, we consider image-prototype-based retrieval, which uses the image prototype only to retrieve the most similar real image from the training set without using it for generation.
@@
Additional ablation studies and extended analyses are provided in Appendix C, including comprehensive analyses of PDS-related hyperparameters, distilled dataset sizes, and alternative decoder conditioning strategies. This appendix also examines robustness on rare samples, evaluates generalization to a different CLIP backbone, and reports results on an additional dataset. Appendix D presents two applications of our method, demonstrating how the distilled dataset can be used within ASIF (Norelli et al., 2023) and how PDS can be incorporated into an optimization-based distillation framework.
@@ C.7 Effect of the distilled dataset size on performance
C.8 Alternative CLIP backbone
For our main experiments, we adopt the ViT-L/14 CLIP model because it is more computationally efficient for extracting image and text features. We additionally evaluate PDS using the ViT-H/14 CLIP to examine whether the improvement generalizes to a different CLIP backbone. As shown in Table 14, using the ViT-H/14 CLIP results in improved performance, likely due to its stronger alignment between image and text features. Importantly, our PDS framework continues to outperform all other baselines even with this CLIP variant, indicating that the observed gains are not specific to ViT-L/14.
Table 14: Results with an alternative CLIP backbone in the 300-pair setting on Flickr30K. We evaluate PDS using the ViT-H/14 CLIP model to examine whether the improvements generalize across different backbones. The ViT-H/14 CLIP achieves higher performance, likely due to its stronger image-text alignment.
Evaluation Model CLIP backbone IR@1 IR@5 IR@10 TR@1 TR@5 TR@10
@@ D.1 Application of the distilled dataset to ASIF
Table 19: Application of the distilled dataset as anchors in ASIF. To evaluate how well a distilled dataset can serve as compact anchors for ASIF, we construct 300 anchors from Flickr30K for each method, including subset selection, other distillation approaches, and PDS. Using a ResNet image encoder and a CLIP text encoder to compute relative representations, we measure retrieval accuracy on Flickr30K. Anchors produced by PDS consistently outperform those obtained from all other baselines, demonstrating that PDS yields more informative and effective anchors for ASIF.
@@ D.2 Integration of PDS into Optimization-based multimodal distillation
We integrate PDS into existing optimization-based multimodal distillation methods by using the PDS-distilled dataset as an initialization, which is known to be crucial for optimization-based distillation. Although replacing real samples used for initialization with the PDS-distilled dataset provides an even more informative starting point, the subsequent optimization process leads to an architecture-dependent distilled dataset, which undermines the cross-architecture generalization achieved by PDS. As shown in Table 20, initializing the optimization-based method with the PDS-distilled dataset tends to improve its performance, but it still underperforms compared to using PDS alone.

### v2

# arXiv:2602.19756v2[cs.CV] 27 Feb 2026

@@ 1 Introduction
A variety of approaches have been proposed to reduce dataset size, including coreset selection (Farahani and Hekmatfar, 2009; Welling, 2009), dataset filtering (Gadre et al., 2023), and pruning methods (Yang et al., 2023; Mahmoud et al., 2024), which have shown competitive performance in the context of CLIP. However, since they rely on representative subsets of the original dataset, they are effective only with sufficiently large reduced datasets and fail once the subset is too small to preserve semantic diversity. In contrast, dataset distillation (Wang et al., 2018) synthesizes new samples that encode the semantics of the original dataset, producing compact yet highly informative datasets. This makes it particularly effective on extremely reduced datasets, where it has outperformed selection-based strategies in image classification (Zhao and Bilen, 2021; Cazenavette et al., 2022), even with only a few distilled samples per class.
@@ 3.1 Problem formulation
where $\mathcal{L}$ denotes a contrastive loss such as InfoNCE. In practice, since text is inherently discrete, dataset distillation is performed in the text embedding space, while images are parameterized in the pixel space. Existing methods typically rely on optimization-based approaches, which require repeated training on the full dataset and storage of intermediate parameters, resulting in significant computational and storage overhead. Moreover, their performance is architecture-dependent, often degrading when applied to models with different backbones.
@@ 3.2 Our PDS framework
Cluster matching for prototype construction. While we obtain semantically coherent clusters within each modality, semantically related image–text clusters are not aligned across modalities. To establish cross-modal correspondences, we measure the association between $C_{i}^{\text{img}}$ and $C_{j}^{\text{txt}}$ by the number of shared image-text pairs whose embeddings belong to the respective clusters. Clusters with more shared pairs are regarded as more semantically related, and we therefore aim to find a one-to-one cluster matching that maximizes the total number of shared pairs.
@@ 4.1 Experimental Settings
Implementation details. Unless otherwise specified, all baselines use CLIP ViT-L/14 (Radford et al., 2021). For PDS, images are generated with the unCLIP decoder using classifier-free guidance (Ho and Salimans, 2022) with a guidance scale of 5.0 and 100 sampling steps, and resized to $224\times 224$ for evaluation. To assess cross-architecture generalization, we follow Xu et al. (2024) and evaluate whether the distilled datasets remain effective when transferred to alternative vision backbones such as ResNet-50 (He et al., 2016) and ViT-Ti/16 (Dosovitskiy et al., 2021). For this evaluation, we use each alternative backbone as the image encoder and attach a learnable linear projection layer after the text encoder. We then fine-tune the CLIP-style model on the distilled dataset while keeping the text encoder frozen. For all baselines, we adopt the hyperparameters reported in the original papers, and conduct all experiments on a single RTX 3090 GPU.
@@ 4.2 Performance Comparison
We evaluate PDS against three categories of baselines: multimodal dataset distillation methods, dataset subset selection methods, and learning-free distillation methods for image classification. These comparisons are designed to highlight three aspects of our contribution: (i) mitigating the cross-generalization limitations of existing multimodal dataset distillation methods, (ii) demonstrating the advantage of PDS over subset selection when the reduced dataset is extremely small, and (iii) overcoming the limitations of existing learning-free distillation approaches when extending from image classification to multimodal datasets. Detailed descriptions of each baseline are provided in Appendix A.
@@
Dataset subset selection. To obtain a highly reduced yet informative dataset, representative images must preserve broad semantic coverage. In this case, dataset distillation offers a more effective solution than subset selection, as it synthesizes data to preserve diverse semantic features. To validate this, we evaluate PDS against classical coreset-selection methods, K-center (Farahani and Hekmatfar, 2009) and Herding (Welling, 2009), as well as widely used filtering schemes such as CLIP score filtering, LAION filtering, and image-based filtering (Gadre et al., 2023).
As shown in Table 2, PDS consistently achieves the highest performance, outperforming the strongest baseline (Herding) by 17.2 and 10.8 $\mathrm{pp}$ in IR@10 and TR@10, respectively. Although dataset filtering methods are widely used in CLIP training due to their effectiveness with moderately reduced datasets, they fail with extremely reduced datasets. These results demonstrate that PDS is particularly well-suited for such scenarios, as it preserves broader semantic coverage by interpolating diverse semantic features and achieves superior performance.
@@ 4.3 Ablation studies
Advantage of using a generative model. We analyze the benefits of using a generative model by comparing PDS with optimization-based multimodal dataset distillation methods. Such optimization-based approaches often synthesize images nearly identical to the initialization sample (Figure 3, left), essentially adding architecture-dependent adversarial perturbations rather than generating genuinely new data, thereby limiting semantic diversity. In contrast, PDS employs a generative model to synthesize images from prototypes constructed by interpolating diverse semantic features.
@@
Impact of image prototype-based image generation. To assess the impact of image prototype-based generation, we compare PDS with three baselines that obtain images through different mechanisms. First, we consider two baselines that do not use image prototypes. The first is text prototype-based retrieval, which retrieves the caption in the training set whose embedding has the highest cosine similarity to the given text prototype and then selects its paired real image. The second is an unCLIP-based text-to-image generation method (Ramesh et al., 2022), which applies the full unCLIP pipeline (prior and decoder) to generate an image from the retrieved caption. In addition, we consider image prototype-based retrieval, which uses the image prototype only to retrieve the most similar real image from the training set without using it for generation.
@@
Additional ablation studies and extended analyses are provided in Appendix C, including comprehensive analyses of PDS-related hyperparameters, distilled dataset sizes, and alternative decoder conditioning strategies. This appendix also examines robustness on rare samples, evaluates generalization to a different CLIP variant, and reports results on an additional dataset. Appendix D presents two applications of our method, demonstrating how the distilled dataset can be used within ASIF (Norelli et al., 2023) and how PDS can be incorporated into an optimization-based distillation framework.
@@ C.7 Effect of the distilled dataset size on performance
C.8 Alternative CLIP variant
There are two publicly available pre-trained unCLIP models based on CLIP ViT-L/14 and ViT-H/14, respectively. For our main experiments, we adopt CLIP ViT-L/14 because it is more computationally efficient for extracting image and text features. We additionally evaluate PDS using CLIP ViT-H/14 to examine whether the improvement generalizes to a different CLIP variant. As shown in Table 14, using CLIP ViT-H/14 results in improved performance, likely due to its stronger alignment between image and text features. Importantly, our PDS framework continues to outperform all other baselines even with this CLIP variant, indicating that the observed gains are not specific to CLIP ViT-L/14.
Table 14: Results with an alternative CLIP variant in the 300-pair setting on Flickr30K. We evaluate PDS using CLIP ViT-H/14 to examine whether the improvements generalize across a different variant. CLIP ViT-H/14 achieves higher performance, likely due to its stronger image-text alignment.
Evaluation Model CLIP variant IR@1 IR@5 IR@10 TR@1 TR@5 TR@10
@@ D.1 Application of the distilled dataset to ASIF
Table 19: Application of the distilled dataset as anchors in ASIF. To evaluate how well distilled or selected datasets can serve as compact anchors for ASIF, we construct 300 anchors from Flickr30K for each method, including subset selection, other distillation approaches, and PDS. Using a ResNet image encoder and a CLIP text encoder to compute relative representations, we measure retrieval accuracy on Flickr30K. Anchors produced by PDS consistently outperform those obtained from all other baselines, demonstrating that PDS yields more informative and effective anchors for ASIF.
@@ D.2 Integration of PDS into Optimization-based multimodal distillation
We integrate PDS into existing optimization-based multimodal distillation methods by using the PDS-distilled dataset as an initialization, which is known to be crucial for optimization-based distillation. Although replacing real samples used for initialization with the PDS-distilled dataset provides a more informative starting point, the subsequent optimization process leads to an architecture-dependent distilled dataset, which undermines the cross-architecture generalization achieved by PDS. As shown in Table 20, initializing the optimization-based method with the PDS-distilled dataset tends to improve its performance, but it still underperforms compared to using PDS alone.